# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v12

@@ 1 Introduction
Machine learning (ML) has progressed rapidly, yet several foundational gaps persist. The field still lacks a unified logical framework that spans the full lifecycle from prior specification and data-driven learning to deployment and online adaptation. Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [1, 2, 3, 4, 5]. Model performance is also hard to predict, which inflates development costs in data, computation, energy, and human effort. In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [6, 7, 8, 9]. Ethical safety lacks a settled mathematical definition and verification protocol. Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [10, 11, 12, 13, 14]. Interpretability similarly lacks a rigorous, domain-agnostic foundation. Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [15, 16, 17, 18, 19].
Information science offers a route to unification [20, 21], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics. Objective Information Theory (OIT) addresses this by axiomatizing information as an enabling mapping from ontological states to carrier states. Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [22, 23]. Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [24]. We complete a remaining gap by formalizing state as an interpretation of well-formed formulas (WFFs) in a higher-order logical system, and by clarifying that enabling mappings are surjective mathematical mappings that are also physically realizable causal relations. We further introduce noisy information and information chains, which support analysis of multi-stage pipelines and their causal dependencies.
@@
• Error control with observable quantities. We derive explicit total variation distance (TVD) upper bounds in two regimes: conservative prediction in regions not covered by training under noise-free input, and arbitrary prediction under noisy input. The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [1, 25, 26, 27, 28, 29, 30, 31, 32].
@@ 2.1 Conceptualizations of Machine Learning
Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [1]. Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [2]. Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [3]. Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [4, 5]. These lines of work clarify individual components of the learning pipeline, yet they do not supply a unified semantic and causal framework that spans model construction, training, input, inference, and deployment.
@@ 2.2 Interpretability
Despite extensive interest, interpretability lacks a precise and generally accepted definition and evaluation protocol. Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [15]. In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [16]. Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [17]. Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [15, 16, 17, 18, 19].
@@ 2.3 Ethical Safety Assurance
As AI systems permeate high-stakes domains, ethical safety has drawn sustained attention in law, policy, and engineering. Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [10, 11]. In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [12]. Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [13]. Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [14]. What remains missing is a mathematically grounded definition of ethical safety with constructible guarantees and verifiable criteria that integrate with ML pipelines.
@@ 2.4 Model Error Estimation
Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [1]. Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [25]. Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [26]. In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [27]. For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [28]. These advances quantify error in specific settings but do not explicitly model causal semantics across multi-stage information flows, nor do they integrate noise structure, coverage geometry, and uncovered-region strategies into a single computable bound that can guide data auditing and prediction strategy design.
@@ 2.5 Toward a Unified Information-Theoretic and Logical Perspective
Information science offers a unifying lens, yet its subfields have developed largely in silos [20, 21]. OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [22, 23, 24]. By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [22, 23, 33, 34]. The present work builds on this foundation to offer a single formal semantics for lifecycle reasoning about interpretability, ethical safety, and computable TVD-based error control.
@@ 3 Theoretical Basis
3.1 Information as Enabling Mapping
[23] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”. Here, we proceed to further clarify the concept.
@@ 3.1 Information as Enabling Mapping
Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [33], we obtain a more robust theoretical foundation for defining information:
@@
The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [34]. Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [35].
3.2 Noisy Information
[22] states that an information instance $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ admits an inverse mapping $I^{-1}(S_{c}(c,T_{m}))$. In practice, however, information realization is rarely immune to noise or other external factors. Consequently, applying $I^{-1}$ may fail to fully recover $S_{o}(o,T_{h})$; instead, it yields a state corrupted by noise. This motivates the following notion.
Definition 3 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, the inverse $I^{-1}(S_{c}(c,T_{m}))$ is another state of the ontology $o$ at time $T_{h}$, denoted by $\widetilde{S_{o}}\left(o,T_{h}\right)$, and given by
@@ 3.2 Noisy Information
where $S_{o}^{-}$ is the noise-loss state with $S_{o}^{-}\subseteq S_{o}$, and $S_{o}^{+}$ is the noise-addition state with $S_{o}{\cap S}_{o}^{+}=\varnothing$. The carrier state $S_{c}(c,T_{m})$ can thus be viewed as the image of $\widetilde{S_{o}}\left(o,T_{h}\right)$ under an enabling mapping:
$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows S_{c}\left(c,T_{m}\right)$ (3.3)
which is the noisy information corresponding to $I$, while $I$ itself is the noise-free information.
Fig. 1(b) depicts the structural characteristics of noisy information. In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [35]. This structural decomposition will support our TVD bounds by isolating loss and addition masses.
@@ 3.3 Information Chain
[22] uses composition to discuss the transitivity of information. Specifically, if $I_{o}:S_{o}(o,T_{h})\rightrightarrows S_{c_{1}}(c_{1},T_{m_{1}})$ and $I_{1}:S_{c_{1}}(c_{1},T_{m_{1}})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ are two information instances, then $I_{2}:S_{o}(o,T_{h})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ is also an information instance, interpreted as $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$. This yields the notion of an information chain.

### v13

@@ 1 Introduction
Machine learning (ML) has progressed rapidly, yet several foundational gaps persist. The field still lacks a unified logical framework that spans the full lifecycle from prior specification and data-driven learning to deployment and online adaptation. Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [vapnik1999overview, alquier2024user, scholkopf2021toward, garcez2023neurosymbolic, marra2024statistical]. Model performance is also hard to predict, which inflates development costs in data, computation, energy, and human effort. In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [achiam2023gpt4, touvron2023llama, chowdhery2023palm, gunasekar2023textbooks]. Ethical safety lacks a settled mathematical definition and verification protocol. Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [russell2015research, burr2023ethical, tilala2024ethical, hawkins2021guidance, pant2024ethics]. Interpretability similarly lacks a rigorous, domain-agnostic foundation. Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [lipton2018mythos, glanois2024survey, zhang2024robustness, rudin2019stop, miller2019explanation].
Information science offers a route to unification [shwartz2017opening, xu2022information], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics. Objective Information Theory (OIT) addresses this by axiomatizing information as an enabling mapping from ontological states to carrier states. Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [xu2014objective, xu2024research]. Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [xu2025general]. We complete a remaining gap by formalizing state as an interpretation of well-formed formulas (WFFs) in a higher-order logical system, and by clarifying that enabling mappings are surjective mathematical mappings that are also physically realizable causal relations. We further introduce noisy information and information chains, which support analysis of multi-stage pipelines and their causal dependencies.
@@
• Error control with observable quantities. We derive explicit total variation distance (TVD) upper bounds in two regimes: conservative prediction in regions not covered by training under noise-free input, and arbitrary prediction under noisy input. The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [vapnik1999overview, bartlett2005local, xu2017information, wang2024theoretical, huang2025convergence, goodfellow2016deep, pinsker1964information, rodriguez2021tighter, devroye2013probabilistic].
@@ 2.1 Conceptualizations of Machine Learning
Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [vapnik1999overview]. Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [alquier2024user]. Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [scholkopf2021toward]. Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [garcez2023neurosymbolic, marra2024statistical]. These lines of work clarify individual components of the learning pipeline, yet they do not supply a unified semantic and causal framework that spans model construction, training, input, inference, and deployment.
@@ 2.2 Interpretability
Despite extensive interest, interpretability lacks a precise and generally accepted definition and evaluation protocol. Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [lipton2018mythos]. In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [glanois2024survey]. Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [zhang2024robustness]. Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [lipton2018mythos, glanois2024survey, zhang2024robustness, rudin2019stop, miller2019explanation].
@@ 2.3 Ethical Safety Assurance
As AI systems permeate high-stakes domains, ethical safety has drawn sustained attention in law, policy, and engineering. Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [russell2015research, burr2023ethical]. In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [tilala2024ethical]. Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [hawkins2021guidance]. Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [pant2024ethics]. What remains missing is a mathematically grounded definition of ethical safety with constructible guarantees and verifiable criteria that integrate with ML pipelines.
@@ 2.4 Model Error Estimation
Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [vapnik1999overview]. Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [bartlett2005local]. Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [xu2017information]. In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [wang2024theoretical]. For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [huang2025convergence]. These advances quantify error in specific settings but do not explicitly model causal semantics across multi-stage information flows, nor do they integrate noise structure, coverage geometry, and uncovered-region strategies into a single computable bound that can guide data auditing and prediction strategy design.
@@ 2.5 Toward a Unified Information-Theoretic and Logical Perspective
Information science offers a unifying lens, yet its subfields have developed largely in silos [shwartz2017opening, xu2022information]. OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [xu2014objective, xu2024research, xu2025general]. By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [xu2014objective, xu2024research, qiu2025research, wiener2019cybernetics]. The present work builds on this foundation to offer a single formal semantics for lifecycle reasoning about interpretability, ethical safety, and computable TVD-based error control.
@@ 3 Theoretical Basis
3.1 A Unified Information Model of Semantic Intension, Statistical Characteristics, and Physical Realization
3.1.1 Information as Enabling Mapping
[xu2024research] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”. Here, we proceed to further clarify the concept.
@@ 3.1.1 Information as Enabling Mapping
Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [qiu2025research], we obtain a more robust theoretical foundation for defining information:
@@
The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [wiener2019cybernetics]. Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [shannon1948mathematical].
3.1.2 Ideal Information
Definition 3 (Semantic Inclusion) Let $S$ be a foundational semantic system. For a logical system $L$, if there exists a semantic mapping $\mathrm{Sem}_{L}$ from $L$ into $S$ such that, for every symbol $x$ in $L$, $x$ is synonymous with $\mathrm{Sem}_{L}(x)$, then $S$ is said to semantically include $L$. If $L_{1}$ and $L_{2}$ are both semantically included by $S$, then $L_{1}$ and $L_{2}$ are said to be synonymous relative to $S$.
Definition 4 (Synonymous States) Let $S_{x_{1}}(x_{1},t_{1})$ and $S_{x_{2}}(x_{2},t_{2})$ be sets of states expressible in logical systems $L_{1}$ and $L_{2}$, respectively. If $L_{1}$ and $L_{2}$ are synonymous relative to $S$, and for every $\varphi_{1}\in S_{x_{1}}(x_{1},t_{1})$ there exists $\varphi_{2}\in S_{x_{2}}(x_{2},t_{2})$, and for every $\varphi_{2}\in S_{x_{2}}(x_{2},t_{2})$ there exists $\varphi_{1}\in S_{x_{1}}(x_{1},t_{1})$, such that $\mathrm{Sem}_{L_{1}}(\varphi_{1})=\mathrm{Sem}_{L_{2}}(\varphi_{2})$, then $S_{x_{1}}(x_{1},t_{1})$ and $S_{x_{2}}(x_{2},t_{2})$ are said to be synonymous relative to $S$, or simply synonymous.
Definition 5 (Ideal Information) Consider information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$. If $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to the foundational semantic system $S$, then $I$ is called ideal information with respect to $S$; when no confusion arises, we may simply call $I$ ideal information.
Since the ontological state $S_{o}(o,T_{h})$ and the carrier state $S_{c}(c,T_{m})$ of ideal information $I$ are synonymous, either $S_{o}(o,T_{h})$ or $S_{c}(c,T_{m})$ can be taken to represent $I$.
3.1.3 The Semantic Intension of Information
@@
Definition 7 (Logical Consequences and Implicational Information Amount) For the ontological state $S_{o}$ of ideal information $I$, the logical closure $\mathrm{Cn}(S_{o})$ contains all logical consequences of $I$, and $|\mathrm{Cn}(S_{o})|$ is the implicational information amount of $I$.
Definition 8 (Atomic Information and Primitive Information Amount) For the ontological state $S_{o}$ of ideal information $I$, suppose $\varphi$ cannot be derived from the other WFFs in $S_{o}$; then $\varphi$ is called an atomic state of $S_{o}$, and $I(\varphi)$ is an atomic piece of information in $I$. Let $\mathrm{Atom}(S_{o})=\{\varphi\mid\varphi\text{ is an atomic state in }S_{o}\}$. Then $|\mathrm{Atom}(S_{o})|$ is the primitive information amount of $I$.
Explicit assertions, logical consequences, and ontic atomic states all belong to the semantic intension of information.
Definition 9 (Logical Depth) For the ontological state $S_{o}$ of ideal information $I$, let $\varphi\in S_{o}$ or $\varphi\in\mathrm{Cn}(S_{o})$. If deriving $\varphi$ from the atomic states of $S_{o}$ requires $i$ ($i=0,1,\dotsm$) recursive layers, then the logical depth of $\varphi$ in $S_{o}$ is $i$, denoted $\mathrm{Ld}(\varphi)$. The logical depths of $S_{o}$ and $\mathrm{Cn}(S_{o})$ are defined by $\mathrm{Ld}(S_{o})=\max\{\mathrm{Ld}(\varphi)\mid\varphi\in S_{o}\}$ and $\mathrm{Ld}(\mathrm{Cn}(S_{o}))=\max\{\mathrm{Ld}(\varphi)\mid\varphi\in\mathrm{Cn}(S_{o})\}$, respectively.
3.1.4 The Extensional Characteristics of Information
Definition 10 (Extensional Characteristics of Information) The extensional characteristics of information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ in a logical system $L$ are the states $S(S_{o}\cup S_{c})$ expressible in $L$ by $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$, denoted $\mathrm{Ext}(I)$.
Statistical characteristics are an important component of extensional characteristics. For ideal information $I$, the explicit information amount $|S_{o}|$, the implicational information amount $|\mathrm{Cn}(S_{o})|$, the primitive information amount $|\mathrm{Atom}(S_{o})|$, and the logical depths $\mathrm{Ld}(S_{o})$ and $\mathrm{Ld}(\mathrm{Cn}(S_{o}))$ are all statistical characteristics and belong to $\mathrm{Ext}(I)$.
Definition 11 (Information Entropy) For the ontological state $S_{o}=\{\varphi_{i}\mid i=1,\dotsm,|S_{o}|\}$ of ideal information $I$, suppose the probability of each $\varphi_{i}$ is $p_{i}$ ($i=1,\dotsm,|S_{o}|$). Then $H(I)=\sum_{i=1}^{|S_{o}|}p_{i}$ is the information entropy of $I$. It also belongs to the statistical characteristics of $\mathrm{Ext}(I)$.
3.1.5 Physical Realization of Information
The physical realization of information is an enabling mapping. For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, practical circumstances—due to constraints such as channel conditions, environment, energy, and time—make it difficult to ensure that $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to the foundational semantic system $S$.
Definition 12 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, it is not difficult to show that there always exists a set $\widetilde{S_{o}}(o,T_{h})=S_{o}(o,T_{h})\setminus S_{o}^{-}(o,T_{h})\cup S_{o}^{+}(o,T_{h})$ such that $\widetilde{S_{o}}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to $S$. Here $S_{o}^{-}$ is called the noise-loss state set, with $S_{o}^{-}\subseteq S_{o}$; $S_{o}^{+}$ is called the noise-addition state set, with $S_{o}\cap S_{o}^{+}=\emptyset$. In this case, $S_{c}(c,T_{m})$ can also be viewed as the result of an enabling mapping from $\widetilde{S_{o}}(o,T_{h})$, namely $\widetilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, which is called the noisy information of $I$.
Fig. 1(b) depicts the structural characteristics of noisy information. In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [shannon1948mathematical]. This structural decomposition will support our TVD bounds by isolating loss and addition masses.
@@ 3.2 Information Chain
[xu2014objective] uses composition to discuss the transitivity of information. Specifically, if $I_{o}:S_{o}(o,T_{h})\rightrightarrows S_{c_{1}}(c_{1},T_{m_{1}})$ and $I_{1}:S_{c_{1}}(c_{1},T_{m_{1}})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ are two information instances, then $I_{2}:S_{o}(o,T_{h})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ is also an information instance, interpreted as $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$. This yields the notion of an information chain.